### v1

# On the existence of magic squares of powers

Abstract.
For any $d\geqslant 2$, we prove that there exists an integer $n_{0}(d)$ such that there exists an $n\times n$ magic square of $d^{\textnormal{th}}$ powers for all $n\geqslant n_{0}(d)$. In particular, we establish the existence of an $n\times n$ magic square of squares for all $n\geqslant 4$, which settles a conjecture of Várilly-Alvarado.
All previous approaches had been based on constructive methods and the existence of $n\times n$ magic squares of $d^{\textnormal{th}}$ powers had only been known for sparse values of $n$. We prove our result by the Hardy-Littlewood circle method, which in this setting essentially reduces the problem to finding a sufficient number of disjoint linearly independent subsets of the columns of the coefficient matrix of the equations defining magic squares. We prove an optimal (up to a constant) lower bound for this quantity. We believe this is the first instance of the circle method being applied to study magic squares.
1. Introduction
Let $n\geqslant 1$ be an integer. A magic square is an $n\times n$ grid of distinct positive integers whose columns, rows, and two major diagonals all sum to the same number. The number to which all rows, columns and diagonals sum is known as the square’s magic constant.
Magic squares have a long and rich history. Legend has it that the earliest recorded $3\times 3$ magic square was first observed by Emperor Yu on the shell of a sacred turtle, which emerged from the waters of the Lo River [Cam60, pp.118]. Since then magic squares have appeared in various cultures, and have been an object of curiosity in art, philosophy, religion and mathematics. The study of magic squares with additional structure is a topic that has garnered great interest in both recreational and research mathematics.
The first $4\times 4$ magic square of squares (Figure 1) was constructed by Euler, in a letter sent to Lagrange in $1770$. Though Euler did not provide any explanation of how he constructed the square, he presented his method to the St. Petersburg Academy of Sciences the same year; the construction is based on the observation that the product of two sums of four squares can itself be expressed as a sum of four squares. This idea was used in $1754$ by Euler to make partial progress, which led Lagrange, in the same year as the letter, to the first complete proof of the four square theorem: every positive integer is the sum of at most four squares [Boy05, VA21].
$\begin{array}[]{|c|c|c|c|}\hline\cr 68^{2}&29^{2}&41^{2}&37^{2}\\ \hline\cr 17^{2}&31^{2}&79^{2}&32^{2}\\ \hline\cr 59^{2}&28^{2}&23^{2}&61^{2}\\ \hline\cr 11^{2}&77^{2}&8^{2}&49^{2}\\ \hline\cr\end{array}$
Figure 1. Euler’s $4\times 4$ magic square of squares with magic constant $8515$.
The search for a $3\times 3$ magic square of squares was popularized by Martin Gardner in $1996$ offering a \$$100$ prize to the first person to construct such a square, though the problem had already been posed by Edouard Lucas in $1876$ and Martin LaBar in $1984$ [Boy05, VA21]. The prize remains unclaimed. However, there are number of results making progress on this problem, which a comprehensive list can be found in [Boy]. A $3\times 3$ magic square of squares gives rise to a rational point with nonzero coordinates on a surface cut out by $6$ quadrics in the space $\mathbb{R}^{8}$. A deep conjecture of Lang predicts that the rational points on this surface are contained within finitely many curves. The curves of genus at least 2 can contain only finitely many rational points by Falting’s theorem. Moreover, the method of [BTVA22] establishes that indeed this surface contains only finitely many curves of genus 00 or $1$. In fact, we know the surface contains curves of genus 00 or $1$ that whose points do not correspond to magic squares. Therefore, it seems plausible that perhaps there are no $3\times 3$ magic squares of squares, or that they are remarkably rare.
An $n\times n$ magic square of squares is a rational point on a variety cut out by $2n$ quadrics in the space $\mathbb{R}^{n^{2}-1}$. In contrast to the $n=3$ case, it is quite reasonable from a geometric point of view that these spaces would carry many rational points for $n\geqslant 5$ [VA21]. This line of logic led Várilly-Alvarado to make the following conjecture.
Conjecture 1.1 ([VA21, Conjecture 4.3]).
There is a positive integer $n_{0}(2)$ such that for every integer $n\geqslant n_{0}(2)$, there exists an $n\times n$ magic square of squares.
In the light of Euler’s example above, Várilly-Alvarado further suggested that the conjecture holds with $n_{0}(2)=4$. We establish that this is indeed the case.
Theorem 1.2.
For every integer $n\geqslant 4$, there exists an $n\times n$ magic square of squares.
In fact, we also establish a generalization of the conjecture for higher powers.
Theorem 1.3.
Let $d\geqslant 3$. There is a positive integer $n_{0}(d)$ such that for every integer $n\geqslant n_{0}(d)$, there exists an $n\times n$ magic square of $d^{\textnormal{th}}$ powers.
The novel feature of our work is that we are applying the Hardy-Littlewood circle method to a problem, for which all previous results had been based on constructive methods. In particular, our result is non-constructive. We believe this is the first instance of the circle method being applied to study magic squares. Prior to our result, the existence of $n\times n$ magic squares of $d^{\textnormal{th}}$ powers was only known for sparse values of $n$ when $d\geqslant 4$ (for $n=q^{d}$ with $q\geqslant 2$ and additionally for small values of $n$ when $4\leqslant d\leqslant 7$), and the existence was also known for $n$ in certain congruence classes when $d\in\{2,3\}$, while we establish the result for all $n\geqslant n_{0}(d)$. We present a more detailed description of the progress on this topic in Section 1.1.
As shown in (2.2), the statement of Theorem 1.3 hold with
$n_{0}(d)=\begin{cases}14\min\{2^{d},d(d+1)\}+79&\mbox{if }3\leqslant d\leqslant 19,\\ 14\lceil d(\log d+4.20032)\rceil+79&\mbox{if }d\geqslant 20.\end{cases}$ (1.1)
It is worth mentioning that the smallest known example of magic squares of $d^{\textnormal{th}}$ powers was of size $2^{d}$ for $d\geqslant 8$; therefore, we improve on the smallest size of a magic square of $d^{\textnormal{th}}$ powers known to exist from $2^{d}$ to $n_{0}(d)$ as in (1.1) for $d>10$.
The main challenge in applying the circle method to the equations defining magic squares is that they define a variety which is “too singular” for the method to be directly applicable; this is explained in Section 1.2. Though we can not apply the results for systems of general homogeneous forms by Birch [Bir62] and Rydin Myerson [RM18], we can apply the version available for systems of diagonal forms by Brüdern and Cook [BC92]. However, this result requires certain “partitionability” of the coefficient matrix, which amounts to finding a sufficient number of disjoint linearly independent subsets of the columns of the coefficient matrix; establishing this is the main technical challenge of the paper. We prove an optimal (up to a constant) lower bound for this quantity in Theorem 2.2, which is the main technical result of this paper. In Appendix A, we turn the proof of Theorem 2.2 into an algorithm and establish a stronger lower bound by explicit computation for small values of $n$. In this context, the circle method allow us to transform the problem of considerable complexity of finding magic squares of $d^{\textnormal{th}}$ powers to a theoretically and computationally simpler problem of finding disjoint linearly independent subsets of the columns of the coefficient matrix. We expect that this lower bound will be particularly useful for studying other variants of magic squares. In fact, in our follow-up work we apply the techniques developed in this paper to study multimagic squares (defined in Section 1.1). We believe that our Theorems 1.2 and 1.3 are a significant step towards the complete classification of the existence of $n\times n$ magic squares of $d^{\textnormal{th}}$ powers.
1.1. Progress on magic squares of powers
An $n\times n$ magic square is a $d$-multimagic square if it remains a magic square when all the entries are raised to the $i^{\text{th}}$ power for every $i=2,\ldots,d$. It is called a normal $d$-multimagic square if it is a $d$-multimagic square with entries consisting from $1$ to $d^{2}$. Clearly a normal $d$-multimagic square provides an example of a magic square of $d^{\text{th}}$ powers. The first published $2$-multimagic squares, which were of sizes $8\times 8$ and $9\times 9$, were obtained by Pfeffermann in $1890$. The first $3$-multimagic square was obtained in $1905$ by Tarry. In fact, Tarry was the first to devise a systematic method of constructing $2$-mulitmagic and $3$-multumagic squares [Kee11]. Prior to our Theorem 1.3, all known examples of magic squares of $d^{\text{th}}$ powers for $d>7$ in fact came from multimagic squares. For $2$-multimagic squares it was proved by Chen, Li, Pan and Xu [PLCX21] that there exists a $2m\times 2m$ normal $2$-multimagic squares for all $m\geqslant 4$. There is also a result by Hu, Meng, Pan, Su and Xiong [HMP+23] which establishes the existence of $16m\times 16m$ normal $3$-multimagic squares for all $m\geqslant 1$. We refer the reader to the introductions of [PLCX21], [HMP+23] and [HP24] for more detailed history on the progress regarding normal $2$ and $3$-multimagic squares. The most general result regarding multimagic squares is given by Derksen, Eggermont and van der Essen [DEvdE07] who have proved that there exist $n\times n$ normal $d$-multimagic squares for any $n=q^{d}$ with $q\geqslant 2$. There is also a comprehensive list of various magic squares recorded by Boyer [Boy] which includes the following:
  • $n\times n$ magic squares of squares for $4\leqslant n\leqslant 7$.
  • $n\times n$ $2$-multimagic squares for $8\leqslant n\leqslant 64$.
  • number of magic squares of $d^{\text{th}}$ powers for $2\leqslant d\leqslant 7$.
Thus we see that our Theorems 1.2 and 1.3 are the first result of this type and of different nature compared to all previous results in this research area.
1.2. Application of the Hardy-Littlewood circle method
Let us label the entries of the $n\times n$ magic square of $d^{\text{th}}$ powers as follows:
$\begin{array}[]{|c|c|c|}\hline\cr x_{1,1}^{d}&\ldots&x_{1,n}^{d}\\ \hline\cr\vdots&\vdots&\vdots\\ \hline\cr x_{n,1}^{d}&\ldots&x_{n,n}^{d}\\ \hline\cr\end{array}\,.$
Let $\mu$ be a positive integer. The system of equations defining the $n\times n$ magic square of $d^{\textnormal{th}}$ powers with magic constant $\mu$ is equivalent to
$x_{1,1}^{d}+\cdots+x_{1,n}^{d}$ $=$ $\mu$
$\vdots$
$x_{n,1}^{d}+\cdots+x_{n,n}^{d}$ $=$ $\mu$
$x_{1,1}^{d}+\cdots+x_{n,1}^{d}$ $=$ $\mu$
$\vdots$
$x_{1,n-1}^{d}+\cdots+x_{n,n-1}^{d}$ $=$ $\mu$
$x_{1,1}^{d}+\cdots+x_{n,n}^{d}$ $=$ $\mu$
$x_{1,n}^{d}+\cdots+x_{n,1}^{d}$ $=$ $\mu,$
which we denote by $\mathbf{F}_{0}(\mathbf{x}_{0})=\bm{\mu}$. A priori the system of equations defining the $n\times n$ magic square of $d^{\textnormal{th}}$ powers with magic constant $\mu$ require $22n+$ 2 equations; however, it can be verified that there is one degree of redundancy (the $12n-$ 1 equations corresponding to $n$ rows and $n-1$ columns imply the equation corresponding to the remaining column). Therefore, this is a system of equations defined by $R_{0}=2n+1$ degree $d$ homogeneous polynomials in $N_{0}=n^{2}$ variables. Since the number of variables $N_{0}$ grows quadratically with $R_{0}$, it may seem reasonable at first to expect that the known results regarding the Hardy-Littlewood circle method readily apply to this system. As it turns out, the system (1.2) is “too singular” for this to be the case.
A seminal result by Birch [Bir62] establishes the existence of a non-trivial integral solution to a general system of equations $F_{1}(\mathbf{x})=\cdots=F_{R}(\mathbf{x})=0$ defined by $R$ homogeneous polynomials of degree $d$ in $N$ variables. For $d=2$, this was greatly improved by Rydin Myerson [RM18]. An important quantity in these results is the dimension of the singular locus of the pencil
$\sigma_{\mathbb{R}}(\mathbf{F})=\max_{\bm{\beta}\in\mathbb{R}^{R}}\dim\{\mathbf{x}_{0}\in\mathbb{A}^{N}:\bm{\beta}.\nabla\mathbf{F}(\mathbf{x})=\mathbf{0}\},$
where $\mathbf{F}=(F_{1},\ldots,F_{R})$. The required bound for $\sigma_{\mathbb{R}}(\mathbf{F})$ is $N>\sigma_{\mathbb{R}}(\mathbf{F})+(d-1)2^{d-1}R(R+1)$ for Birch’s result to be applicable, while $N>\sigma_{\mathbb{R}}(\mathbf{F})+8R$ for Rydin Myerson’s result. However, it can be verified that
$\sigma_{\mathbb{R}}(\mathbf{F}_{0})\geqslant n^{2}-n=N_{0}-\frac{R_{0}-1}{2},$
which is far too large for either of the mentioned results to be applicable. Both of these results are for a general system of homogeneous forms; however, the system (1.2) consists only of diagonal polynomials. There are results in this direction as well. Let us now suppose that $F_{1},\ldots,F_{R}$ are diagonal polynomials. The system of diagonal equations $F_{1}(\mathbf{x})=\cdots=F_{R}(\mathbf{x})=0$ was first studied by Davenport and Lewis [DL69, Lemma 32]; their result required
$N\geqslant\begin{cases}\lfloor 9R^{2}d\log(3Rd)\rfloor&\mbox{if $d$ is odd},\\ \lfloor 48R^{2}d^{3}\log(3Rd^{2})\rfloor&\mbox{if $d\geqslant 4$ is even},\end{cases}$
to establish the existence of a non-trivial integral solution. Since $N_{0}=(R_{0}-1)^{2}/4$, we can not hope to directly apply this result to (1.2). By incorporating the breakthrough on Waring’s problem by Vaughan [Vau89], Brüdern and Cook [BC92] improved the number of variables required to be
$N\geqslant 2dR(\log d+O(\log\log d)).$
We remark that both of these results require a suitable “rank condition” on the coefficient matrix. Since the time of these two papers, there has been great progress regarding Waring’s problem (for example, by Wooley [Woo92], [Woo95] and more recently by Wooley and Brüdern [BW23]) and also the resolution of Vinogradov’s mean value theorem (see the work by Bourgain, Demeter and Guth [BDG16], and by Wooley [Woo16, Woo19]). By incorporating these recent developments, the required number of variables may be further improved. Though this technical procedure is mostly standard, we present the details in our separate paper [RS] and the result applied to (1.2) is summarised in Theorem 2.2.
Acknowledgements
NR was supported by FWF project ESP 441-NBL while SY by a FWF grant (DOI 10.55776/P32428). The authors are grateful to Tim Browning for his constant encouragement and enthusiasm, and Jörg Brüdern and Trevor Wooley for very helpful discussions regarding his paper [BC92] and recent developments in Waring’s problem, respectively, and Diyuan Wu for turning the proof of Theorem 2.2 into an algorithm and running the computation, as explained in Appendix A. They would also like to thank Christian Boyer for maintaining his website [Boy] which contains a comprehensive list of various magic squares discovered, Brady Haran and Tony Várilly-Alvarado for their public engagement activity of mathematics and magic squares of squares [^1] A popular YouTube video “Magic Squares of Squares (are PROBABLY impossible)” of the Numberphile channel by Brady Haran, in which Tony Várilly-Alvarado appears as a guest speaker: https://www.youtube.com/watch?v=Kdsj84UdeYg., and all the magic squares enthusiasts who have contributed to [Boy], which made this paper possible.
Notation
We make use of the standard abbreviations $e(z)=e^{2\pi iz}$ and $e_{q}(z)=e^{\frac{2\pi iz}{q}}$. Given a vector $\mathbf{a}=(a_{1},\ldots,a_{R})\in\mathbb{Z}^{R}$, by $1\leqslant\mathbf{a}\leqslant q$ we mean $1\leqslant a_{i}\leqslant R$ for each $1\leqslant i\leqslant R$.
2. Magic squares of powers
Let $\mu$ be a positive integer. Let $M_{0}$ be the coefficient matrix of the system (1.2), which is given by (5.1). We define
=$\mathscr{I}=\{0\}\cup\{(i_{1},j_{1};i_{2},j_{2}):(i_{1},j_{1})\neq(i_{2},j_{2}),1\leqslant i_{1},j_{1},i_{2},j_{2}\leqslant d\}.$
For each $\sigma=(i_{1},j_{1};i_{2},j_{2})\in\mathscr{I}\setminus\{0\}$, we denote $\mathbf{x}_{\sigma}$ to be $\mathbf{x}_{0}$ with $x_{i_{2},j_{2}}$ removed and $\mathbf{F}_{\sigma}(\mathbf{x}_{\sigma})=\bm{\mu}$ the system of equations obtained by substituting $x_{i_{2},j_{2}}=x_{i_{1},j_{1}}$ into $\mathbf{F}_{0}(\mathbf{x}_{0})=\bm{\mu}$. We then denote by $M_{\sigma}$ the coefficient matrix of this system, which is obtained from $M_{0}$ by adding the $x_{i_{2},j_{2}}$-column to the $x_{i_{1},j_{1}}$-column then deleting the $x_{i_{2},j_{2}}$-column.
For $\sigma\in\mathscr{I}$, $\mathfrak{B}\subseteq\mathbb{N}$ and $X\geqslant 1$, we introduce the following counting functions
$N_{\sigma}(\mathfrak{B};X,\mu)=\#\{\mathbf{x}_{\sigma}\in(\mathfrak{B}\cap[1,X])^{n^{2}-\epsilon(\sigma)}:\mathbf{F}_{\sigma}(\mathbf{x}_{\sigma})=\bm{\mu}\},$
where
$\epsilon(\sigma)=\begin{cases}0&\mbox{if }\sigma=0,\\ 1&\mbox{if }\sigma\in\mathscr{I}\setminus\{0\}.\end{cases}$
Then the number of magic squares, whose entries are restricted to $\mathfrak{B}$, with magic constant $\mu$ is then given by
$N(\mathfrak{B};X,\mu)=N_{0}(\mathfrak{B};X,\mu)+O\left(\sum_{\sigma\in\mathscr{I}\setminus\{0\}}N_{\sigma}(\mathfrak{B};X,\mu)\right).$ (2.1)
In this expression, the sum in the $O$-term is the contribution from squares whose entries are not distinct. We estimate these counting functions using the Hardy-Littlewood circle method for $\mathfrak{B}=\mathbb{N}$ and
$\mathscr{A}(X,X^{\eta})=\{x\in[1,X]\cap\mathbb{N}:\textnormal{prime }p|x\textnormal{ implies }p\leqslant X^{\eta}\}$
for sufficiently small $\eta>0$.
Definition 2.1.
Let $\sigma\in\mathscr{I}$. We define $\Psi({M_{\sigma}})$ to be the largest integer $T$ such that there exists
$\{\mathfrak{D}_{1},\ldots,\mathfrak{D}_{T}\},$
where each $\mathfrak{D}_{i}$ is a linearly independent set of $12n+$ 1 columns of $M_{\sigma}$ and $\mathfrak{D}_{i}\cap\mathfrak{D}_{j}\neq\varnothing$ if $i\neq j$. We let $T_{\sigma}$ be a non-negative integer such that
$\Psi({M_{\sigma}})\geqslant T_{\sigma}.$
By applying the Hardy-Littlewood method to the system of diagonal equations (1.2), we obtain the following as a direct consequence of [RS, Theorems 1 and 2].
Theorem 2.2.
Let $\mathfrak{B}=\mathbb{N}$ or $\mathscr{A}(X,X^{\eta})$ with $\eta>0$ sufficiently small. Suppose
$\min_{\sigma\in\mathscr{I}}T_{\sigma}\geqslant\begin{cases}\min\{2^{d}+1,d(d+1)+1\}&\mbox{if }2\leqslant d<19\mbox{ and }\mathfrak{B}=\mathbb{N},\\ \lceil d(\log d+4.20032)\rceil+1&\mbox{if }d\geqslant 20\mbox{ and }\mathfrak{B}=\mathscr{A}(X,X^{\eta}).\end{cases}$
Then there exists $\lambda>0$ such that
$N_{0}(\mathfrak{B};X,\mu)=C_{\mathfrak{B}}\mathfrak{S}\mathfrak{I}X^{n^{2}-(2n+1)d}+O(X^{n^{2}-(2n+1)d}(\log X)^{-\lambda}),$
where
$C_{\mathfrak{B}}=\begin{cases}1&\mbox{if }\mathfrak{B}=\mathbb{N},\\ c(\eta)&\mbox{if }\mathfrak{B}=\mathscr{A}(X,X^{\eta}),\end{cases}$
$c(\eta)>0$ is a constant depending only on $\eta$, the singular series $\mathfrak{S}$ and the singular integral $\mathfrak{I}$ are defined in (3.1) and (3.3), respectively. Furthermore,
$\max_{\sigma\in\mathscr{I}\setminus\{0\}}N_{\sigma}(\mathfrak{B};X,\mu)\ll X^{n^{2}-(2n+1)d-1}.$
The next step is to establish lower bounds for $\mathfrak{S}$ and $\mathfrak{I}$ which are independent of $X$. We present the details of the following proposition in Section 3.
Proposition 2.3.
Given any $X\geqslant 1$ sufficiently large, there exists $\mu=\mu(X)\in\mathbb{N}$ such that
$\mathfrak{S}\mathfrak{I}>\mathfrak{c}_{n,d},$
where $\mathfrak{c}_{n,d}>0$ is a constant depending only on $n$ and $d$.
Finally, in order to establish the existence of magic squares of $d^{\textnormal{th}}$ powers with magic constant $\mu$, it remains to establish the existence of $T_{\sigma}$. The following is the main technical result of the paper, which we prove through Sections 47.
Theorem 2.4.
We have
$T_{0}\geqslant\begin{cases}\frac{n-28}{12}&\mbox{if $n$ is even,}\\ \frac{n-37}{14}&\mbox{if $n$ is odd.}\end{cases}$
By combining these results collected in this section, we obtain Theorem 1.3 as an immediate consequence.
Proof of Theorem 1.3.
Since
$\min_{\sigma\in\mathscr{I}\setminus\{0\}}T_{\sigma}\geqslant T_{0}-2,$
it follows from Theorem 2.4 that
$\min_{\sigma\in\mathscr{I}}T_{\sigma}\geqslant\begin{cases}\frac{n-28}{12}-2&\mbox{if $n$ is even,}\\ \frac{n-37}{14}-2&\mbox{if $n$ is odd.}\end{cases}$
Therefore, by combining this estimate, Theorem 2.2 and Proposition 2.3 with (2.1), we obtain that Theorem 1.3 holds with
$n_{0}(d)=\begin{cases}12\min\{2^{d},d(d+1)\}+64&\mbox{if $2\leqslant d\leqslant 19$ and $n$ is even,}\\ 12\lceil d(\log d+4.20032)\rceil+64&\mbox{if $d\geqslant 20$ and $n$ is even,}\\ 14\min\{2^{d},d(d+1)\}+79&\mbox{if $2\leqslant d\leqslant 19$ and $n$ is odd,}\\ 14\lceil d(\log d+4.20032)\rceil+79&\mbox{if $d\geqslant 20$ and $n$ is odd.}\end{cases}$ (2.2)
In order to establish Theorem 1.2 we need to deal with smaller values of $n$. In Appendix A, we will turn the proof of Theorem 2.4 into an algorithm and establish a lower bound for $\min_{\sigma\in\mathscr{I}}T_{\sigma}$ computationally for these $n$.
Proof of Theorem 1.2.
It follows from the proof of Theorem 1.3 above that there exists an $n\times n$ magic square of squares as soon as
$n\geqslant\begin{cases}112&\mbox{if $n$ is even,}\\ 135&\mbox{if $n$ is odd.}\end{cases}$
It is easy to see from Table 1 in Appendix A that
$\min_{\sigma\in\mathscr{I}}T_{\sigma}\geqslant 5$
for $50\leqslant n\leqslant 140$. Thus for these values of $n$ the existence of $n\times n$ magic squares of squares follows by Theorem 2.2. Finally, since the statement is already known for $4\leqslant n\leqslant 64$, in fact explicit examples have been discovered and listed in [Boy], this completes the proof. ∎
3. Singular series and singular integral: Proof of Proposition 2.3
We let $\textnormal{Col}(M_{0})$ denote the set of columns of $M_{0}$. Let $\textnormal{Jac}_{\mathbf{F}_{0}}$ denote the Jacobian matrix of $\mathbf{F}_{0}$ and $\mathbb{F}=\mathbb{R}$ or $\mathbb{Q}_{p}$ for any prime $p$. A crucial fact we make use of in this section is that given any $z\in\mathbb{F}\setminus\{0\}$, we have
$\textnormal{Jac}_{\mathbf{F}_{0}}(z,\ldots,z)=dz^{d-1}M_{0},$
which in particular is of full rank. Therefore, if $\mathbf{F}_{0}(z,\ldots,z)=\bm{\mu}$, then it is in fact a non-singular solution.
Let
$S(q,a)=\sum_{1\leqslant x\leqslant q}e_{q}(ax^{d}).$
We define the singular series
$\mathfrak{S}=\sum_{q=1}^{\infty}A(q),$ (3.1)
where
$A(q)=q^{-n^{2}}\sum_{\begin{subarray}{c}1\leqslant\mathbf{a}\leqslant q\\ \gcd(q,\mathbf{a})=1\end{subarray}}\prod_{\mathbf{c}\in\textnormal{Col}(M)}S(q,\mathbf{a}.\mathbf{c})\cdot e_{q}\left(-\mu\sum_{i=1}^{2n+1}a_{i}\right).$
We have the following regarding $\mathfrak{S}$ which is [RS, Lemmas 3.1 and 3.2].
Lemma 3.1.
Suppose
$T_{0}>\frac{d(2n+2)}{2n+1}.$
Then
$A(q)\ll q^{-(2n+1)(\frac{T_{0}}{d}-1)}$ (3.2)
and the series (2.3) converges absolutely. In fact,
$\mathfrak{S}=\prod_{p\textnormal{ prime}}\chi(p),$
where
$\chi(p)=1+\sum_{k=1}^{\infty}A(p^{k}).$
We establish the desired lower bound for $\mathfrak{S}$ for a special value of $\mu$.
Lemma 3.2.
Let $\mu=np_{0}^{d}$, where $p_{0}$ is a prime sufficiently large with respect to $n$ and $d$. Then $\mathfrak{S}_{0}>c_{n,d}$, where $c_{n,d}>0$ is a constant depending only on $n$ and $d$.
Proof.
An application of (3.2) yields
$|\chi(p)-1|\ll\sum_{k=1}^{\infty}p^{-k(2n+1)(\frac{T_{0}}{d}-1)}\ll p^{-(2n+1)(\frac{T_{0}}{d}-1)}.$
Therefore, there exists $P>0$ such that
$\prod_{p>P}\chi(p)>\frac{1}{2}.$
We remark that $P$ is independent of $\mu$. The standard argument show that
$\chi(p)=\lim_{m\rightarrow\infty}\frac{\nu(p^{m})}{p^{m(n^{2}-2n-1)}},$
where $\nu(p^{m})$ is the number of solutions to (1.2) in $\left(\mathbb{Z}/p^{m}\mathbb{Z}\right)^{n^{2}}$. Let $\nu_{p}(d)$ denote the $p$-adic valuation of $d$. Since $\mu=np_{0}^{d}$ with prime $p_{0}$ sufficiently large, it follows that there exists a non-singular solution to $\mathbf{F}_{0}(\mathbf{x}_{0})\equiv\bm{\mu}\,(\operatorname{mod}{p^{\nu_{p}(d)+1}})$ for every prime $p\leqslant P$. It then follows by Hensel’s lemma that
$\chi(p)=\lim_{m\rightarrow\infty}\frac{\nu(p^{m})}{p^{m(n^{2}-2n-1)}}\geqslant\lim_{m\rightarrow\infty}\frac{p^{(m-(\nu_{p}(d)+1))(n^{2}-2n-1)}}{p^{m(n^{2}-2n-1)}}=p^{-(\nu_{p}(d)+1)(n^{2}-2n-1)}.$
Therefore, we obtain
$\mathfrak{S}>\frac{1}{2}\prod_{p\leqslant P}p^{-(\nu_{p}(d)+1)(n^{2}-2n-1)}\gg 1,$
where the implicit constant is independent of $\mu$. ∎
Let
$I(\beta)=\int_{0}^{1}e(\beta\xi^{d})\mathrm{d}\xi.$
We define the singular integral
$\mathfrak{I}=\int_{\mathbb{R}^{2n+1}}\prod_{\mathbf{c}\in\textnormal{Col}(M_{0})}I(\bm{\gamma}.\mathbf{c})\cdot e\left(-\frac{\mu}{X^{d}}\sum_{i=1}^{2n+1}\gamma_{i}\right)\mathrm{d}\bm{\gamma}.$ (3.3)
We have the following regarding $\mathfrak{I}$ which is [RS, Lemmas 3.3]
Lemma 3.3.
Suppose $T_{0}>d$. Then the integral (3.3) converges absolutely.
We establish the desired lower bound for $\mathfrak{I}$ for a special value of $\mu$.
Lemma 3.4.
Let $\varepsilon_{0}>0$ be sufficiently small. Suppose $\mu=n\zeta^{d}X^{d}$ with $\zeta\in[\varepsilon_{0},1-\varepsilon_{0}]$. Then $\mathfrak{I}_{0}>c_{n,d,\varepsilon_{0}}$, where $c_{n,d,\varepsilon_{0}}$ is a positive constant depending only on $n,d$ and $\varepsilon_{0}$.
Proof.
It follows by the standard argument as in [Sch82] that
$\mathfrak{I}=\lim_{L\rightarrow\infty}\int_{[0,1]^{n^{2}}}\prod_{i=1}^{2n+1}\Phi_{L}(F_{0,i}(\mathbf{x}_{0})-\mu X^{-d})\mathrm{d}\mathbf{x}_{0},$
where
$\Phi_{L}(\eta)=\begin{cases}L(1-L|\eta|)&\mbox{if }|\eta|\leqslant L^{-1},\\ 0&\mbox{otherwise. }\end{cases}$
Furthermore, since $\mu X^{-d}=n\zeta^{d}$ with $\zeta\in[\varepsilon_{0},1-\varepsilon_{0}]$, we see that $\mathbf{x}_{0}=(\zeta,\ldots,\zeta)$ is a non-singular solution to the system of equations $\mathbf{F}_{0}(\mathbf{x}_{0})=\bm{\mu}$. It then follows by an application of the implicit function theorem as in [Sch82, Lemma 2] that $\mathfrak{I}>0$. In particular, since $\zeta\in[\varepsilon_{0},1-\varepsilon_{0}]$, the lower bound is independent of $\mu$. ∎
Let $\varepsilon_{0}>0$ be sufficiently small. By the prime number theorem, for any $X$ sufficiently large there exists a prime $p_{0}$ satisfying
$\varepsilon_{0}X<p_{0}<(1-\varepsilon_{0})X.$
Then for $\mu=np_{0}^{d}$ we see that Lemmas 3.2 and 3.4 hold, from which Proposition 2.3 follows.
4. Linear equations
Lemma 4.1.
The system of $2k$ linear equations in $12k+$ 1 variables
$x_{1}+y_{1}=0$ $x_{2}+y_{1}=0$
$\vdots$
$x_{k-1}+y_{k-1}=0$ $x_{k}+y_{k-1}=0$
$x_{k}+y_{k}=0$ $z+y_{k}=0$
is equivalent to
$z=x_{1}=\cdots=x_{k}=-y_{1}=\cdots=-y_{k}.$
Proof.
It follows from the $12k-$ 1 equations without $z+y_{k}=0$ that
$x_{1}=\cdots=x_{k}=-y_{1}=\cdots=-y_{k}.$
Then it is immediate that
$z=-y_{i}=x_{i}$
holds for all $1\leqslant i\leqslant k$ as desired. ∎
Let $\mathbb{L}_{n}$ be the following system of $2n$ linear equations
$x_{1}+y_{1}=0$ $x_{2}+y_{1}=0$
$x_{2}+y_{2}=0$ $x_{3}+y_{2}=0$
$\vdots$
$x_{n-1}+y_{n-1}=0$ $x_{n}+y_{n-1}=0$
$x_{n}+y_{n}=0$ $x_{1}+y_{n}=0.$
We denote by $\mathbb{L}_{n}(m;i_{1},k_{1};i_{2},k_{2})$ the system of $2n$ linear equations obtained by replacing the equations
$x_{i_{1}+1}+y_{i_{1}}=0,\,x_{k_{1}+1}+y_{k_{1}}=0,\,x_{i_{2}+1}+y_{i_{2}}=0,\,x_{k_{2}+1}+y_{k_{2}}=0,$
in $\mathbb{L}_{n}$ (where we identify $x_{n+1}$ as $x_{1}$) with
$x_{i_{1}+1}+y_{i_{1}}+x_{k_{1}+1}=0,\,y_{k_{1}}=0,\,x_{i_{2}+1}+y_{i_{2}}+x_{k_{2}+1}=0,\,y_{k_{2}}=0,$
respectively, and then replacing the equation
$x_{m+1}+y_{m}=0$
with $y_{m}=0$.
Given $1\leqslant s,t\leqslant n$, we shall use the following notation
$[s,t]=\begin{cases}\{s,s+1,\ldots,t\}&\mbox{if }s\leqslant t,\\ \{s,s+1,\ldots,n,1,2,\ldots,t\}&\mbox{if }s>t.\end{cases}$
Lemma 4.2.
Suppose $1\leqslant i_{1}<k_{1}<i_{2}<k_{2}\leqslant n$ holds and $m\in[k_{1}+1,i_{2}-1]$. Then
$\{(\mathbf{x},\mathbf{y})\in\mathbb{R}^{2n}:(\mathbf{x},\mathbf{y})\textnormal{ satisfies }\mathbb{L}_{n}(m;i_{1},k_{1};i_{2},k_{2})\}=\{\mathbf{0}\}.$
Proof.
Consider the system of equations $\mathbb{L}_{n}(m;i_{1},k_{1};i_{2},k_{2})$. Since $y_{k_{2}}=0$, it follows that $x_{k_{2}}=0$. Then it follows that $y_{k_{2}-1}=0$ and consequently $x_{k_{2}-1}=0$. This pattern continues until we reach $y_{i_{2}+1}=x_{i_{2}+1}=0$, which implies $x_{k_{2}+1}+y_{i_{2}}=0$. Therefore, we have
$x_{i}=y_{i}=0\quad(i\in[i_{2}+1,k_{2}])$
and the remaining equations are
$x_{k_{2}+1}+y_{k_{2}+1}=0$ $x_{k_{2}+2}+y_{k_{2}+1}=0$
$x_{k_{2}+2}+y_{k_{2}+2}=0$ $x_{k_{2}+3}+y_{k_{2}+2}=0$
$\vdots$
$x_{i_{1}-1}+y_{i_{1}-1}=0$ $x_{i_{1}}+y_{i_{1}-1}=0$
$x_{i_{1}}+y_{i_{1}}=0$ $x_{i_{1}+1}+y_{i_{1}}+x_{k_{1}+1}=0$
$x_{i_{1}+1}+y_{i_{1}+1}=0$ $x_{i_{1}+2}+y_{i_{1}+1}=0$
$\vdots$
$x_{k_{1}-1}+y_{k_{1}-1}=0$ $x_{k_{1}}+y_{k_{1}-1}=0$
$x_{k_{1}}+y_{k_{1}}=0$ $y_{k_{1}}=0$
$x_{k_{1}+1}+y_{k_{1}+1}=0$ $x_{k_{1}+2}+y_{k_{1}+1}=0$
$\vdots$
$x_{i_{2}-1}+y_{i_{2}-1}=0$ $x_{i_{2}}+y_{i_{2}-1}=0$
$x_{i_{2}}+y_{i_{2}}=0$ $x_{k_{2}+1}+y_{i_{2}}=0,$
with the understanding that the equation $x_{m+1}+y_{m}=0$ is replaced with $y_{m}=0$. We note that we have written the system of equations as above for convenience using the symmetry of the equations (and also by identifying $x_{n+r}$ and $y_{n+r}$ as $x_{r}$ and $y_{r}$ respectively).
We now repeat the argument starting from $y_{k_{1}}=0$ until we obtain $y_{i_{1}+1}=x_{i_{1}+1}=0$. Then we have
$x_{i}=y_{i}=0\quad(i\in[i_{1}+1,k_{1}]\cup[i_{2}+1,k_{2}])$
and the remaining equations are
$x_{k_{2}+1}+y_{k_{2}+1}=0$ $x_{k_{2}+2}+y_{k_{2}+1}=0$
$x_{k_{2}+2}+y_{k_{2}+2}=0$ $x_{k_{2}+3}+y_{k_{2}+2}=0$
$\vdots$
$x_{i_{1}-1}+y_{i_{1}-1}=0$ $x_{i_{1}}+y_{i_{1}-1}=0$
$x_{i_{1}}+y_{i_{1}}=0$ $x_{k_{1}+1}+y_{i_{1}}=0$
$x_{k_{1}+1}+y_{k_{1}+1}=0$ $x_{k_{1}+2}+y_{k_{1}+1}=0$
$\vdots$
$x_{i_{2}-1}+y_{i_{2}-1}=0$ $x_{i_{2}}+y_{i_{2}-1}=0$
$x_{i_{2}}+y_{i_{2}}=0$ $x_{k_{2}+1}+y_{i_{2}}=0.$
By Lemma 4.1 the system of equations
$x_{k_{2}+1}+y_{k_{2}+1}=0$ $x_{k_{2}+2}+y_{k_{2}+1}=0$
$\vdots$
$x_{i_{1}-1}+y_{i_{1}-1}=0$ $x_{i_{1}}+y_{i_{1}-1}=0$
$x_{i_{1}}+y_{i_{1}}=0$ $x_{k_{1}+1}+y_{i_{1}}=0$
is equivalent to
$x_{i}=-y_{i}=x_{k_{1}+1}$ (4.1)
for all $i\in[k_{2}+1,i_{1}]$. Similarly, the system of equations
$x_{k_{1}+1}+y_{k_{1}+1}=0$ $x_{k_{1}+2}+y_{k_{1}+1}=0$
$\vdots$
$x_{m-1}+y_{m-1}=0$ $x_{m}+y_{m-1}=0$
$x_{m}+y_{m}=0$ $y_{m}=0$
$x_{m+1}+y_{m+1}=0$ $x_{m+2}+y_{m+1}=0$
$\vdots$
$x_{i_{2}-1}+y_{i_{2}-1}=0$ $x_{i_{2}}+y_{i_{2}-1}=0$
$x_{i_{2}}+y_{i_{2}}=0$ $x_{k_{2}+1}+y_{i_{2}}=0$
is equivalent to
$x_{k_{1}+1}=y_{k_{1}+1}=\cdots=x_{m}=y_{m}=0$ (4.2)
and
$x_{i}=-y_{i}=x_{k_{2}+1}$ (4.3)
for all $i\in[m+1,i_{2}]$. Combining (4.1), (4.2) and (4.3) yields
$x_{i}=y_{i}=0\quad([k_{1}+1,i_{2}]\cup[k_{2}+1,i_{1}]).$
Finally, on recalling that all other $x_{i}$ and $y_{i}$ are 00, the result follows. ∎
Next we denote by $\widetilde{\mathbb{L}}_{n}(m;i_{1},k_{1};i_{2},k_{2})$ the system of $2n$ linear equations obtained by replacing the equations
$x_{i_{1}}+y_{i_{1}}=0,\,x_{k_{1}}+y_{k_{1}}=0,\,x_{i_{2}+1}+y_{i_{2}}=0,\,x_{k_{2}+1}+y_{k_{2}}=0,$
in $\mathbb{L}_{n}$ with
$x_{i_{1}}+y_{i_{1}}+y_{k_{1}}=0,\,x_{k_{1}}=0,\,x_{i_{2}+1}+y_{i_{2}}+x_{k_{2}+1}=0,\,y_{k_{2}}=0,$
respectively, and then replacing the equation
$x_{m+1}+y_{m}=0$
with $y_{m}=0$.
Lemma 4.3.
Suppose $1\leqslant i_{1}<k_{1}<i_{2}<k_{2}\leqslant n$ holds and $m\in[k_{1}+1,i_{2}-1]$. Then
$\{(\mathbf{x},\mathbf{y})\in\mathbb{R}^{2n}:(\mathbf{x},\mathbf{y})\textnormal{ satisfies }\widetilde{\mathbb{L}}_{n}(m;i_{1},k_{1};i_{2},k_{2})\}=\{\mathbf{0}\}.$
Proof.
Consider the system of equations $\widetilde{\mathbb{L}}_{n}(m;i_{1},k_{1};i_{2},k_{2})$. Since $y_{k_{2}}=0$, it follows that $x_{k_{2}}=0$. Then it follows that $y_{k_{2}-1}=0$ and consequently $x_{k_{2}-1}=0$. This pattern continues until we reach $y_{i_{2}+1}=x_{i_{2}+1}=0$, which implies $x_{k_{2}+1}+y_{i_{2}}=0$. Therefore, we have
$x_{i}=y_{i}=0\quad(i\in[i_{2}+1,k_{2}])$
and the remaining equations are
$x_{k_{2}+1}+y_{k_{2}+1}=0$ $x_{k_{2}+2}+y_{k_{2}+1}=0$
$x_{k_{2}+2}+y_{k_{2}+2}=0$ $x_{k_{2}+3}+y_{k_{2}+2}=0$
$\vdots$
$x_{i_{1}-1}+y_{i_{1}-1}=0$ $x_{i_{1}}+y_{i_{1}-1}=0$
$x_{i_{1}}+y_{i_{1}}+y_{k_{1}}=0$ $x_{i_{1}+1}+y_{i_{1}}=0$
$x_{i_{1}+1}+y_{i_{1}+1}=0$ $x_{i_{1}+2}+y_{i_{1}+1}=0$
$\vdots$
$x_{k_{1}-1}+y_{k_{1}-1}=0$ $x_{k_{1}}+y_{k_{1}-1}=0$
$x_{k_{1}}=0$ $x_{k_{1}+1}+y_{k_{1}}=0$
$x_{k_{1}+1}+y_{k_{1}+1}=0$ $x_{k_{1}+2}+y_{k_{1}+1}=0$
$\vdots$
$x_{i_{2}-1}+y_{i_{2}-1}=0$ $x_{i_{2}}+y_{i_{2}-1}=0$
$x_{i_{2}}+y_{i_{2}}=0$ $x_{k_{2}+1}+y_{i_{2}}=0,$
with the understanding that the equation $x_{m+1}+y_{m}=0$ is replaced with $y_{m}=0$.
We now start from $x_{k_{1}}=0$, which implies $y_{k_{1}-1}=0$ and consequently $x_{k_{1}-1}=0$. This pattern continues until we obtain $x_{i_{1}+1}=y_{i_{1}+1}=0$, which finally implies $y_{i_{1}}=0$. Then we have
$x_{k_{1}}=y_{i_{1}}=x_{i}=y_{i}=0\quad(i\in[i_{1}+1,k_{1}-1]\cup[i_{2}+1,k_{2}])$
and the remaining equations are
$x_{k_{2}+1}+y_{k_{2}+1}=0$ $x_{k_{2}+2}+y_{k_{2}+1}=0$
$x_{k_{2}+2}+y_{k_{2}+2}=0$ $x_{k_{2}+3}+y_{k_{2}+2}=0$
$\vdots$
$x_{i_{1}-1}+y_{i_{1}-1}=0$ $x_{i_{1}}+y_{i_{1}-1}=0$
$x_{i_{1}}+y_{k_{1}}=0$
$x_{k_{1}+1}+y_{k_{1}}=0$
$x_{k_{1}+1}+y_{k_{1}+1}=0$ $x_{k_{1}+2}+y_{k_{1}+1}=0$
$\vdots$
$x_{i_{2}-1}+y_{i_{2}-1}=0$ $x_{i_{2}}+y_{i_{2}-1}=0$
$x_{i_{2}}+y_{i_{2}}=0$ $x_{k_{2}+1}+y_{i_{2}}=0.$
It follows from an application of Lemma 4.1 that the system of equations
$x_{k_{2}+1}+y_{k_{2}+1}=0$ $x_{k_{2}+2}+y_{k_{2}+1}=0$
$x_{k_{2}+2}+y_{k_{2}+2}=0$ $x_{k_{2}+3}+y_{k_{2}+2}=0$
$\vdots$
$x_{i_{1}-1}+y_{i_{1}-1}=0$ $x_{i_{1}}+y_{i_{1}-1}=0$
$x_{i_{1}}+y_{k_{1}}=0$
is equivalent to
$x_{i_{1}}=x_{i}=-y_{i}=-y_{k_{1}}$
for all $i\in[k_{2}+1,i_{1}-1]$. Similarly, the system of equations
$x_{k_{1}+1}+y_{k_{1}}=0$
$x_{k_{1}+1}+y_{k_{1}+1}=0$ $x_{k_{1}+2}+y_{k_{1}+1}=0$
$\vdots$
$x_{m-1}+y_{m-1}=0$ $x_{m}+y_{m-1}=0$
$x_{m}+y_{m}=0$ $y_{m}=0$
$x_{m+1}+y_{m+1}=0$ $x_{m+2}+y_{m+1}=0$
$\vdots$
$x_{i_{2}-1}+y_{i_{2}-1}=0$ $x_{i_{2}}+y_{i_{2}-1}=0$
$x_{i_{2}}+y_{i_{2}}=0$ $x_{k_{2}+1}+y_{i_{2}}=0$
is equivalent to
$y_{k_{1}}=x_{k_{1}+1}=y_{k_{1}+1}=\cdots=x_{m}=y_{m}=0$
and
$x_{i}=-y_{i}=x_{k_{2}+1}$
for all $i\in[m+1,i_{2}]$. Thus we have obtained that
$x_{i_{1}}=y_{k_{1}}=x_{1}=y_{i}=0\quad(i\in[k_{1}+1,i_{2}]\cup[k_{2}+1,i_{1}-1]).$
Finally, on recalling that all other $x_{i}$ and $y_{i}$ are 00, the result follows. ∎
5. Preliminaries
Let $I_{n}$ denote the $n\times n$ identity matrix and define $\widetilde{I}_{n}$ to be the $(n-1)\times n$ matrix obtained by removing the last row from $I_{n}$. Then the $(2n+1)\times n^{2}$ coefficient matrix of the system (1.2) is
$M_{0}=\begin{bmatrix}\begin{bmatrix}1&1&\cdots&1\\ 0&0&\cdots&0\\ \vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&0\end{bmatrix}&\begin{bmatrix}0&0&\cdots&0\\ 1&1&\cdots&1\\ \vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&0\end{bmatrix}&\cdots&\begin{bmatrix}0&0&\cdots&0\\ 0&0&\cdots&0\\ \vdots&\vdots&\vdots&\vdots\\ 1&1&\cdots&1\end{bmatrix}\\ \widetilde{I}_{n}&\widetilde{I}_{n}&\cdots&\widetilde{I}_{n}\\ \begin{bmatrix}1&0&\cdots&0\\ 0&0&\cdots&1\end{bmatrix}&\begin{bmatrix}0&1&\cdots&0&0\\ 0&0&\cdots&1&0\end{bmatrix}&\cdots&\begin{bmatrix}0&0&\cdots&1\\ 1&0&\cdots&0\end{bmatrix}\end{bmatrix}.$ (5.1)
We shall denote by $\mathbf{c}_{i,j}$ the $((i-1)n+j)$-th column of $M_{0}$, for each $1\leqslant i,j\leqslant n$. The first subscript $i$ will always satisfy $1\leqslant i\leqslant n$, but for the second subscript $j$, to simplify our exposition, we will consider it modulo $n$, that is
$c_{i,j+mn}=c_{i,j},$
for any $1\leqslant j\leqslant n$ and $m\in\mathbb{Z}$.
Definition 5.1.
We shall refer to vectors of the form $c_{i,*}$ by vectors from the $i$-th block and $c_{*,j}$ by vectors from the $j$-th column.
Let
$\mathscr{S}=\mathscr{S}_{1}\cup\mathscr{S}_{2},$
where
$\mathscr{S}_{1}=\{\mathbf{c}_{1,1},\mathbf{c}_{2,2},\ldots,\mathbf{c}_{n-1,n-1},\mathbf{c}_{n,n}\}$ (5.2)
and
$\mathscr{S}_{2}=\{\mathbf{c}_{1,n},\mathbf{c}_{2,n-1},\ldots,\mathbf{c}_{n-1,2},\mathbf{c}_{n,1}\}\setminus\mathscr{S}_{1}.$ (5.3)
The columns in $\mathscr{S}_{1}$ are precisely the ones with $1$ in the $2n$-th entry and $\mathscr{S}_{2}$ the ones with $1$ in the $(2n+1)$-th entry that are not contained in $\mathscr{S}_{1}$.
Let $1\leqslant N_{1}<N_{2}\leqslant n$ to be chosen in due course. In particular, we choose it so that
$N_{2}<\frac{n}{2}-1.$ (5.4)
For each $N_{1}\leqslant\ell\leqslant N_{2}$, we define
$\widetilde{\mathfrak{B}}_{\ell}=\{\mathbf{c}_{1,1+(2\ell-1)},\mathbf{c}_{2,2+(2\ell-1)},\ldots,\mathbf{c}_{n,n+(2\ell-1)}\}\bigcup\left(\{\mathbf{c}_{1,1+2\ell},\mathbf{c}_{2,2+2\ell},\ldots,\mathbf{c}_{n,n+2\ell}\}\setminus\{\mathbf{c}_{n-2\ell,n}\}\right).$
This set contains precisely one vector from the $n$-th column $\mathbf{c}_{n-(2\ell-1),n}$ whose entries between $(n+1)$-th and $(2n-1)$-th position are all 00.
By (5.4) it follows that
$\widetilde{\mathfrak{B}}_{\ell}\cap\widetilde{\mathfrak{B}}_{\ell^{\prime}}=\varnothing$ (5.5)
if $\ell\neq\ell^{\prime}$.
Lemma 5.2.
Given any $N_{1}\leqslant\ell\leqslant N_{2}$, we have
$\widetilde{\mathfrak{B}}_{\ell}\cap\mathscr{S}=\left\{\mathbf{c}_{\frac{n-\epsilon(n)}{2}+1-\ell,\frac{n-\epsilon(n)}{2}+\epsilon(n)+\ell},\mathbf{c}_{n+1-\ell,\ell}\right\},$
where
$\epsilon(n)=\begin{cases}0&\mbox{if $n$ is even},\\ 1&\mbox{if $n$ is odd}.\end{cases}$
Proof.
First it is clear from the definition of $\widetilde{\mathfrak{B}}_{\ell}$ and (5.4) that
$\widetilde{\mathfrak{B}}_{\ell}\cap\mathscr{S}_{1}=\varnothing.$
Therefore, if $\mathbf{c}_{i,j}\in\widetilde{\mathfrak{B}}_{\ell}\cap\mathscr{S}$ with $1\leqslant i,j\leqslant n$, then there exists $\epsilon\in\{0,1\}$ such that
$(i,i+2\ell-1+\epsilon+m(i,\ell,\epsilon)n)=(j,n+1-j),$
where
$m(i,\ell,\epsilon)=\begin{cases}0&\mbox{if }1\leqslant i+2\ell-1+\epsilon\leqslant n,\\ -1&\mbox{if }n<i+2\ell-1+\epsilon\leqslant 2n.\end{cases}$
It follows that $i=j$ and
$j+2\ell-1+\epsilon+m(j,\ell,\epsilon)n=n+1-j,$
or equivalently
$j=\frac{n-\epsilon}{2}+1-\ell-\frac{m(j,\ell,\epsilon)n}{2}.$
Suppose $m(j,\ell,\epsilon)=0$. Then it must be that $\epsilon=0$ if $n$ is even and $1$ if $n$ is odd. Thus we see that
$\mathbf{c}_{\frac{n-\epsilon}{2}+1-\ell,\frac{n-\epsilon}{2}+\epsilon+\ell}\in\widetilde{\mathfrak{B}}_{\ell}\cap\mathscr{S};$
here we used the fact that $1\leqslant\frac{n+\epsilon}{2}+\ell<n$, which follows from (5.4).
On the other hand, suppose $m(j,\ell,\epsilon)=-1$. Then
$j=n-\frac{\epsilon}{2}+1-\ell.$
In particular, it must be that $\epsilon=0$. Thus we see that
$\mathbf{c}_{n+1-\ell,\ell}\in\widetilde{\mathfrak{B}}_{\ell}\cap\mathscr{S}.$
Finally, we have $\mathbf{c}_{\frac{n-\epsilon}{2}+1-\ell,\frac{n-\epsilon}{2}+\epsilon+\ell}\neq\mathbf{c}_{n+1-\ell,\ell}$, since
$\frac{n-\epsilon}{2}+1-\ell\leqslant\frac{n}{2}<n+1-\ell.$
As a consequence of this lemma, there are precisely two vectors in $\widetilde{\mathfrak{B}}_{\ell}$ whose last two entries are not 00. We now want to replace these two vectors from $\widetilde{\mathfrak{B}}_{\ell}$ with appropriate vectors whose last two entries are 00. We shall consider the two cases when $n$ is even and odd separately in Sections 6 and 7 respectively.
6. Proof of Theorem 2.4: $n$ is even
Given $N_{1}\leqslant\ell\leqslant N_{2}$, the two vectors in $\widetilde{\mathfrak{B}}_{\ell}$ (from Lemma 5.2) we need to replace are
$\mathbf{c}_{\frac{n}{2}+1-\ell,\frac{n}{2}+\ell}=\mathbf{c}_{\frac{n}{2}+1-\ell,\frac{n}{2}+1-\ell+(2\ell-1)}\quad\textnormal{and}\quad\mathbf{c}_{n+1-\ell,\ell}=\mathbf{c}_{n+1-\ell,n+1-\ell+(2\ell-1)};$
we replace $\mathbf{c}_{\frac{n}{2}+1-\ell,\frac{n}{2}+\ell}$ and $\mathbf{c}_{n+1-\ell,\ell}$ with appropriate vectors of the form $\mathbf{c}_{\frac{n}{2}+1-\ell,j_{1}(\ell)}$ and $\mathbf{c}_{n+1-\ell,j_{2}(\ell)}$, respectively, that are not in $\mathscr{S}$. We define
$\mathfrak{B}_{\ell}=(\widetilde{\mathfrak{B}}_{\ell}\setminus\{\mathbf{c}_{\frac{n}{2}+1-\ell,\frac{n}{2}+\ell},\mathbf{c}_{n+1-\ell,\ell}\})\bigcup\{\mathbf{c}_{\frac{n}{2}+1-\ell,j_{1}(\ell)},\mathbf{c}_{n+1-\ell,j_{2}(\ell)}\},$ (6.1)
where $j_{1}(\ell)$ and $j_{2}(\ell)$ are to be chosen in due course. In particular, these values are chosen so that every vector in $\mathfrak{B}_{\ell}$ has 00 for the last two entries, i.e. not in $\mathscr{S}$. It is also clear that $\mathfrak{B}_{\ell}$ contains at most two vectors from each block. We further ensure that
$2\ell<j_{1}(\ell)<n,\quad j_{1}(\ell)\neq\frac{n}{2}+\ell,\frac{n}{2}+\ell+1\quad\textnormal{and}\quad 1\leqslant j_{2}(\ell)<\ell.$ (6.2)
Suppose we have the following linear combination of vectors in $\mathfrak{B}_{\ell}$
$\mathbf{0}$ $=$ $\sum_{\begin{subarray}{c}1\leqslant i\leqslant n\\ i\neq\frac{n}{2}+1-\ell,n+1-\ell\end{subarray}}(x_{i}\mathbf{c}_{i,i+(2\ell-1)}+y_{i}\mathbf{c}_{i,i+2\ell})$
$+x_{\frac{n}{2}+1-\ell}\mathbf{c}_{\frac{n}{2}+1-\ell,j_{1}(\ell)}+y_{\frac{n}{2}+1-\ell}\mathbf{c}_{\frac{n}{2}+1-\ell,\frac{n}{2}+1+\ell}$
$+x_{n+1-\ell}\mathbf{c}_{n+1-\ell,j_{2}(\ell)}+y_{n+1-\ell}\mathbf{c}_{n+1-\ell,n+1+\ell},$
with
$y_{n-2\ell}=0.$
Our goal is to show that $x_{i}=y_{i}=0$ for all $1\leqslant i\leqslant n$. Recall from the definition of $\mathbf{c}_{i,j}\notin\mathscr{S}$ that there is $1$ in the $i$-th and $(n+j)$-th entry, provided $1\leqslant i\leqslant n$ and $11\leqslant j\leqslant n-$ 1, and every other entry is 00. There are precisely three summands that correspond to the vector of the form $\mathbf{c}_{*,j_{1}(\ell)}$; they are
$x_{j_{1}(\ell)-2\ell+1}\mathbf{c}_{j_{1}(\ell)-2\ell+1,j_{1}(\ell)-2\ell+1+(2\ell-1)},\quad y_{j_{1}(\ell)-2\ell}\mathbf{c}_{j_{1}(\ell)-2\ell,j_{1}(\ell)-2\ell+2\ell}\quad\textnormal{and}\quad x_{\frac{n}{2}+1-\ell}\mathbf{c}_{\frac{n}{2}+1-\ell,j_{1}(\ell)}.$
For this, we need
$j_{1}(\ell)-2\ell+1,j_{1}(\ell)-2\ell\notin\left\{\frac{n}{2}+1-\ell,n+1-\ell\right\}$
and
$j_{1}(\ell)\notin\left\{j_{2}(\ell),\frac{n}{2}+1+\ell,n+1+\ell\right\},$
which follow from (6.2). Similarly, there are precisely three summands that correspond to the vector of the form $\mathbf{c}_{*,j_{2}(\ell)}$; they are
$x_{n+j_{2}(\ell)-2\ell+1}\mathbf{c}_{n+j_{2}(\ell)-2\ell+1,n+j_{2}(\ell)-2\ell+1+(2\ell-1)},\quad y_{n+j_{2}(\ell)-2\ell}\mathbf{c}_{n+j_{2}(\ell)-2\ell,n+j_{2}(\ell)-2\ell+2\ell}$
and
$x_{n+1-\ell}\mathbf{c}_{n+1-\ell,j_{2}(\ell)}.$
There is only one vector (other than $y_{n-2\ell}\mathbf{c}_{n-2\ell,n}=\mathbf{0}$) that corresponds to a summand of the form $\mathbf{c}_{*,n}$, which is $x_{n-2\ell+1}\mathbf{c}_{n-2\ell+1,n}$. Since $1<n-2\ell+1<n$, the vector $\mathbf{c}_{n-2\ell+1,n}$ has $1$ in the $(n-2\ell+1)$-th entry and 00 otherwise. Thus the vector equation (6) is equivalent to the following system of linear equations
$x_{i}+y_{i}$ $=$ $0\quad(1\leqslant i\leqslant n)$
$x_{i+1}+y_{i}$ $=$ $0\quad\left(i\neq\frac{n}{2}-\ell,n-\ell,j_{1}(\ell)-2\ell,n+j_{2}(\ell)-2\ell,n-2\ell\right)$
$x_{j_{1}(\ell)-2\ell+1}+y_{j_{1}(\ell)-2\ell}+x_{\frac{n}{2}-\ell+1}$ $=$ 00 0
$y_{\frac{n}{2}-\ell}$ $=$ 00 0
$x_{n+j_{2}(\ell)-2\ell+1}+y_{n+j_{2}(\ell)-2\ell}+x_{n-\ell+1}$ $=$ 00 0
$y_{n-\ell}$ $=$ 00 0
$y_{n-2\ell}$ $=$ $0$
This system is precisely $\mathbb{L}_{n}\left(n-2\ell;j_{1}(\ell)-2\ell,\frac{n}{2}-\ell;n+j_{2}(\ell)-2\ell,n-\ell\right)$. It is worth mentioning that the first set of equations $x_{i}+y_{i}=0$ $(1\leqslant i\leqslant n)$ comes from comparing the entries of the vectors between $1$-st and $n$-th position, while the rest comes from the entries between $(n+1)$-th and $(2n-1)$-th position.
It can be verified that (5.4) and (6.2) ensure
$1\leqslant j_{1}(\ell)-2\ell,\frac{n}{2}-\ell,n+j_{2}(\ell)-2\ell,n-\ell,n-2\ell<n.$
Therefore, to make use of Lemma 4.2 we first choose $j_{1}(\ell)$ and $j_{2}(\ell)$ such that
$1\leqslant j_{1}(\ell)-2\ell<\frac{n}{2}-\ell<n+j_{2}(\ell)-2\ell<n-\ell,$
which in turn follows from
$2N_{2}+1\leqslant j_{1}(\ell)<\frac{n}{2}+N_{1}\quad\textnormal{and}\quad 1\leqslant j_{2}(\ell)<N_{1}.$ (6.4)
We note that the number of choices for $j_{1}(\ell)$ and $j_{2}(\ell)$ satisfying these conditions are
$\frac{n}{2}+N_{1}-2N_{2}-1\quad\textnormal{and}\quad N_{1}-1,$ (6.5)
respectively. It is also clear that the conditions (6.4) imply (6.2). Finally, to apply Lemma 4.2 we further need to ensure
$n-2\ell\in\left[\frac{n}{2}-\ell+1,n+j_{2}(\ell)-2\ell-1\right];$
this condition follows from
$\ell\leqslant\frac{n}{2}-1\quad\textnormal{and}\quad 1\leqslant j_{2}(\ell).$
In particular, both of these conditions are satisfied for all
$N_{1}\leqslant\ell\leqslant N_{2}<\frac{n}{2}-1$
under the assumption that (6.4) holds. Therefore, as long as (5.4) and (6.4) are satisfied, it follows from Lemma 4.1 that
$x_{i}=y_{i}=0$
for all $1\leqslant i\leqslant n$, which on recalling (6) implies that the vectors of $\mathfrak{B}_{\ell}$ form a linearly independent set.
We shall prove that we may choose $j_{1}(\ell)$ and $j_{2}(\ell)$, for each $N_{1}\leqslant\ell\leqslant N_{2}$, such that
$\mathfrak{B}_{\ell}\cap\mathscr{S}=\varnothing,\quad\dim\textnormal{Span}_{\mathbb{R}}\mathfrak{B}_{\ell}=2n-1\quad\textnormal{and}\quad\mathfrak{B}_{\ell^{\prime}}\cap\mathfrak{B}_{\ell}=\varnothing,$ (6.6)
if $N_{1}\leqslant\ell^{\prime}<\ell$. To this end, we define
$\mathfrak{R}_{\ell}=\mathscr{S}\cup\mathfrak{B}_{N_{1}}\cup\mathfrak{B}_{N_{1}+1}\cup\cdots\cup\mathfrak{B}_{\ell-1}\cup\widetilde{\mathfrak{B}}_{\ell}\cup\cdots\cup\widetilde{\mathfrak{B}}_{N_{2}},$
for each $N_{1}\leqslant\ell\leqslant N_{2}$. Every $\widetilde{\mathfrak{B}}_{\ell^{\prime}}$ contains at most two vectors from each block (recall this notion from Definition 5.1), and the same holds for $\mathfrak{B}_{\ell^{\prime}}$, as mentioned in the sentence following (6.1). Thus $\mathfrak{R}_{\ell}$ contains
$\leqslant(2(N_{2}-N_{1}+1)+2)$ (6.7)
vectors from each block. We shall assume that the following hold
$\frac{n}{2}+N_{1}-2N_{2}-1\geqslant N_{1}-1\quad\textnormal{and}\quad N_{1}-2(N_{2}-N_{1})\geqslant 6.$ (6.8)
We note that, as explained above, by choosing $j_{1}(\ell)$ and $j_{2}(\ell)$ to satisfy (6.4) we have $\dim\textnormal{Span}_{\mathbb{R}}\mathfrak{B}_{\ell}=2n-1$. We now show by induction that we may choose $j_{1}(\ell)$ and $j_{2}(\ell)$ so that
$(\mathfrak{B}_{\ell}\setminus\widetilde{\mathfrak{B}}_{\ell})\cap\mathfrak{R}_{\ell}=\varnothing\quad\textnormal{and}\quad\dim\textnormal{Span}_{\mathbb{R}}\mathfrak{B}_{\ell}=2n-1$ (6.9)
hold for each $N_{1}\leqslant\ell\leqslant N_{2}$. We now record the following as a lemma.
Lemma 6.1.
Suppose (6.9) holds. Then we have (6.6).
Proof.
Since we already have $\dim\textnormal{Span}_{\mathbb{R}}\mathfrak{B}_{\ell}=2n-1$ for all $N_{1}\leqslant\ell\leqslant N_{2}$, it suffices to establish the other two relations. First we have from the construction that
$(\mathfrak{B}_{\ell}\cap\widetilde{\mathfrak{B}}_{\ell})\cap\mathscr{S}=\varnothing\quad(N_{1}\leqslant\ell\leqslant N_{2}).$
Then since (6.9) implies
$(\mathfrak{B}_{\ell}\setminus\widetilde{\mathfrak{B}}_{\ell})\cap\mathscr{S}=\varnothing\quad(N_{1}\leqslant\ell\leqslant N_{2}),$
we have
$\mathfrak{B}_{\ell}\cap\mathscr{S}=\varnothing\quad(N_{1}\leqslant\ell\leqslant N_{2}).$
Next we recall from (5.5) that
$\widetilde{\mathfrak{B}}_{\ell}\cap\widetilde{\mathfrak{B}}_{\ell^{\prime}}=\varnothing$
if $\ell\neq\ell^{\prime}$. Let $N_{1}\leqslant\ell^{\prime}<\ell\leqslant N_{2}$. Then
$\mathfrak{B}_{\ell^{\prime}}\cap\mathfrak{B}_{\ell}$ $=$ $\mathfrak{B}_{\ell^{\prime}}\bigcap(\mathfrak{B}_{\ell}\setminus\widetilde{\mathfrak{B}}_{\ell})\cup(\mathfrak{B}_{\ell}\cap\widetilde{\mathfrak{B}}_{\ell})$
$\subseteq$ $\mathfrak{B}_{\ell^{\prime}}\cap(\mathfrak{B}_{\ell}\setminus\widetilde{\mathfrak{B}}_{\ell})\bigcup\mathfrak{B}_{\ell^{\prime}}\cap\widetilde{\mathfrak{B}}_{\ell}$
$\subseteq$ $\mathfrak{R}_{\ell}\cap(\mathfrak{B}_{\ell}\setminus\widetilde{\mathfrak{B}}_{\ell})\bigcap(\mathfrak{B}_{\ell^{\prime}}\setminus\widetilde{\mathfrak{B}}_{\ell^{\prime}})\cap\widetilde{\mathfrak{B}}_{\ell}$
$\subseteq$ $\mathfrak{R}_{\ell}\cap(\mathfrak{B}_{\ell}\setminus\widetilde{\mathfrak{B}}_{\ell})\bigcap(\mathfrak{B}_{\ell^{\prime}}\setminus\widetilde{\mathfrak{B}}_{\ell^{\prime}})\cap\widetilde{\mathfrak{R}}_{\ell^{\prime}}$
$=$ $\varnothing$,∅ ,
where the final equality is due to (6.9). Thus we have established that (6.6) holds. ∎
We begin with the base case $\ell=N_{1}$. In this case, we know from (6.5), (6.7) and (6.8) that there are
$\geqslant N_{1}-1-(2(N_{2}-N_{1}+1)+2)\geqslant 1$
choices for both $j_{1}(N_{1})$ and $j_{2}(N_{1})$. To see this, first we recall that $j_{1}(N_{1})$ corresponds to the vector $\mathbf{c}_{\frac{n}{2}+1-N_{1},j_{1}(N_{1})}$; we omit the explanation for $j_{2}(N_{1})$ as it is similar. We may select any of the $N_{1}-1$ vectors satisfying (6.4) from the $(\frac{n}{2}+1-N_{1})$-th block which does not belong to $\mathfrak{R}_{N_{1}}$. Since there are $\leqslant(2(N_{2}-N_{1}+1)+2)$ vectors in $\mathfrak{R}_{N_{1}}$ from the $(\frac{n}{2}+1-N_{1})$-th block, we obtain the above inequality. Therefore, we may select appropriate $j_{1}(N_{1})$ and $j_{2}(N_{1})$ such that
$(\mathfrak{B}_{N_{1}}\setminus\widetilde{\mathfrak{B}}_{N_{1}})\cap\mathfrak{R}_{N_{1}}=\varnothing\quad\textnormal{and}\quad\dim\textnormal{Span}_{\mathbb{R}}\mathfrak{B}_{N_{1}}=2n-1,$
and fix these choices.
Now suppose the statement holds for all values greater than or equal to $N_{1}$ and less than $\ell$, for some $N_{1}<\ell\leqslant N_{2}$. We need to choose $j_{1}(\ell)$ and $j_{2}(\ell)$ such that the desired conditions are satisfied, where we recall that these correspond to the vectors $\mathbf{c}_{\frac{n}{2}+1-\ell,j_{1}(\ell)}$ and $\mathbf{c}_{n+1-\ell,j_{2}(\ell)}$ respectively. These two vectors are to be chosen from the complement of $\mathfrak{R}_{\ell}$. Therefore, by the same argument as in the base case above, we again have
$\geqslant N_{1}-1-(2(N_{2}-N_{1}+1)+2)\geqslant 1$
choices for both $j_{1}(\ell)$ and $j_{2}(\ell)$. Thus we may select appropriate $j_{1}(\ell)$ and $j_{2}(\ell)$ such that
$(\mathfrak{B}_{\ell}\setminus\widetilde{\mathfrak{B}}_{\ell})\cap\mathfrak{R}_{\ell}=\varnothing,\quad\textnormal{and}\quad\dim\textnormal{Span}_{\mathbb{R}}\mathfrak{B}_{\ell}=2n-1$
holds. This completes our induction. In conclusion, we have obtained pairwise disjoint $\mathfrak{B}_{N_{1}},\ldots,\mathfrak{B}_{N_{2}}$ satisfying
$\mathfrak{B}_{\ell}\cap\mathscr{S}=\varnothing\quad\textnormal{and}\quad\dim\textnormal{Span}_{\mathbb{R}}\mathfrak{B}_{\ell}=2n-1,$
for each $N_{1}\leqslant\ell\leqslant N_{2}$, provided (5.4) and (6.8) hold.
6.1. Final bound for $T$
Since every vector in $\mathfrak{B}_{\ell}$ has 00 for the last two entries, we obtain the following as an immediate consequence.
Corollary 6.2.
Let $N_{1}\leqslant\ell\leqslant N_{2}$ and $(\mathbf{a},\mathbf{b})\in\mathscr{S}_{1}\times\mathscr{S}_{2}$. Then the vectors of $\mathfrak{B}_{\ell}\cup\{\mathbf{a},\mathbf{b}\}$ form a linearly independent set.
Since
#$\#\mathscr{S}_{1},\#\mathscr{S}_{2}\geqslant n-1\geqslant N_{2}-N_{1}+1$
and $\mathfrak{B}_{\ell}$ are pairwise disjoint, we easily obtain $N_{2}-N_{1}+1$ pairwise disjoint sets of the form
$\mathfrak{D}_{\ell}=\mathfrak{B}_{\ell}\cup\{\mathbf{a}_{\ell},\mathbf{b}_{\ell}\},$
where $(\mathbf{a}_{\ell},\mathbf{b}_{\ell})\in\mathscr{S}_{1}\times\mathscr{S}_{2}$. Thus we have established
$T\geqslant N_{2}-N_{1}+1,$
where $T$ is from the statement of the theorem. Since the only restrictions we needed were (5.4) and (6.8), we set
$N_{2}=\left\lfloor\frac{n}{4}\right\rfloor$
so that
$\frac{n}{2}+N_{1}-2N_{2}-1\geqslant N_{1}-1\quad\textnormal{and}\quad N_{2}<\frac{n}{2}-1$
are satisfied. Then we choose $N_{1}$ in a way that $N_{1}-2(N_{2}-N_{1})\geqslant 6$ holds, for which
$N_{1}=\left\lceil\frac{2}{3}N_{2}+2\right\rceil$
suffices. Then the above bound for $T$ becomes
$T\geqslant N_{2}-\frac{2}{3}N_{2}-2\geqslant\frac{1}{3}\left(\frac{n}{4}-1\right)-2=\frac{n-28}{12}.$
7. Proof of Theorem 2.4: $n$ is odd
The argument in this section is carried out in a similar manner as in Section 6, but there are enough differences that it requires proper explanation. Given $N_{1}\leqslant\ell\leqslant N_{2}$, the two vectors in $\widetilde{\mathfrak{B}}_{\ell}$ (from Lemma 5.2) we need to replace are
$\mathbf{c}_{\frac{n-1}{2}+1-\ell,\frac{n-1}{2}+1+\ell}=\mathbf{c}_{\frac{n-1}{2}+1-\ell,\frac{n-1}{2}+1-\ell+2\ell}\quad\textnormal{and}\quad\mathbf{c}_{n+1-\ell,\ell}=\mathbf{c}_{n+1-\ell,n+1-\ell+(2\ell-1)};$
we replace $\mathbf{c}_{\frac{n-1}{2}+1-\ell,\frac{n-1}{2}+1+\ell}$ and $\mathbf{c}_{n+1-\ell,\ell}$ with appropriate vectors of the form $\mathbf{c}_{i_{1}(\ell),\frac{n+1}{2}+\ell}$ and $\mathbf{c}_{n+1-\ell,j_{2}(\ell)}$, respectively, that are not in $\mathscr{S}$. We define
$\mathfrak{B}_{\ell}=(\widetilde{\mathfrak{B}}_{\ell}\setminus\{\mathbf{c}_{\frac{n-1}{2}+1-\ell,\frac{n-1}{2}+1+\ell},\mathbf{c}_{n+1-\ell,\ell}\})\bigcup\{\mathbf{c}_{i_{1}(\ell),\frac{n-1}{2}+1+\ell},\mathbf{c}_{n+1-\ell,j_{2}(\ell)}\},$
where $i_{1}(\ell)$ and $j_{2}(\ell)$ are to be chosen in due course. In particular, these values are chosen so that every vector in $\mathfrak{B}_{\ell}$ has 00 for the last two entries, i.e. not in $\mathscr{S}$. We further ensure that
$1\displaystyle 1\leqslant i_{1}(\ell)<n+1-\ell,\quad i_{1}(\ell)\neq\frac{n-1}{2}-\ell+1,n-2\ell+$ 1 (7.1)
and
$1\leqslant j_{2}(\ell)<\ell.$
Suppose we have the following linear combination of vectors in $\mathfrak{B}_{\ell}$
$\mathbf{0}$ $=$ $\sum_{\begin{subarray}{c}1\leqslant i\leqslant n\\ i\neq\frac{n-1}{2}+1-\ell,n+1-\ell\end{subarray}}(x_{i}\mathbf{c}_{i,i+(2\ell-1)}+y_{i}\mathbf{c}_{i,i+2\ell})$
$+x_{\frac{n-1}{2}+1-\ell}\mathbf{c}_{\frac{n-1}{2}+1-\ell,\frac{n-1}{2}+\ell}+y_{\frac{n-1}{2}+1-\ell}\mathbf{c}_{i_{1}(\ell),\frac{n-1}{2}+1+\ell}$
$+x_{n+1-\ell}\mathbf{c}_{n+1-\ell,j_{2}(\ell)}+y_{n+1-\ell}\mathbf{c}_{n+1-\ell,n+1+\ell},$
with
$y_{n-2\ell}=0.$
We wish to show $x_{i}=y_{i}=0$ for all $1\leqslant i\leqslant n$. Recall from the definition of $\mathbf{c}_{i,j}\notin\mathscr{S}$ that there is $1$ in the $i$-th and $(n+j)$-th entry, provided $1\leqslant i\leqslant n$ and $11\leqslant j\leqslant n-$ 1, and the remaining entries are all 00. There are precisely three summands that correspond to the vector of the form $\mathbf{c}_{i_{1}(\ell),*}$; they are
$x_{i_{1}(\ell)}\mathbf{c}_{i_{1}(\ell),i_{1}(\ell)+(2\ell-1)},\quad y_{i_{1}(\ell)}\mathbf{c}_{i_{1}(\ell),i_{1}(\ell)+2\ell}\quad\textnormal{and}\quad y_{\frac{n-1}{2}+1-\ell}\mathbf{c}_{i_{1}(\ell),\frac{n-1}{2}+1+\ell}.$
Similarly, there are precisely three summands that correspond to the vector of the form $\mathbf{c}_{*,j_{2}(\ell)}$; they are
$x_{n+j_{2}(\ell)-2\ell+1}\mathbf{c}_{n+j_{2}(\ell)-2\ell+1,n+j_{2}(\ell)-2\ell+1+(2\ell-1)},\quad y_{n+j_{2}(\ell)-2\ell}\mathbf{c}_{n+j_{2}(\ell)-2\ell,n+j_{2}(\ell)-2\ell+2\ell}$
and
$x_{n+1-\ell}\mathbf{c}_{n+1-\ell,j_{2}(\ell)}.$
There is only one vector (other than $y_{n-2\ell}\mathbf{c}_{n-2\ell,n}=\mathbf{0}$) that corresponds to a summand of the form $\mathbf{c}_{*,n}$, which is $x_{n-2\ell+1}\mathbf{c}_{n-2\ell+1,n}$. Since $1<n-2\ell+1<n$, the vector $\mathbf{c}_{n-2\ell+1,n}$ has $1$ in the $(n-2\ell+1)$-th entry and 00 otherwise. Thus the vector equation (7) is equivalent to the following system of linear equations
$x_{i}+y_{i}$ $=$ $0\quad\left(i\neq\frac{n-1}{2}+1-\ell,i_{1}(\ell)\right)$
$x_{i_{1}(\ell)}+y_{i_{1}(\ell)}+y_{\frac{n-1}{2}+1-\ell}$ $=$ 00 0
$x_{\frac{n-1}{2}+1-\ell}$ $=$ 00 0
$x_{i+1}+y_{i}$ $=$ $0\quad\left(i\neq n-\ell,n+j_{2}(\ell)-2\ell,n-2\ell\right)$
$x_{n+j_{2}(\ell)-2\ell+1}+y_{n+j_{2}(\ell)-2\ell}+x_{n+1-\ell}$ $=$ 00 0
$y_{n-\ell}$ $=$ 00 0
$y_{n-2\ell}$ $=$ $0$
This system is precisely $\widetilde{\mathbb{L}}_{n}\left(n-2\ell;i_{1}(\ell),\frac{n-1}{2}+1-\ell;n+j_{2}(\ell)-2\ell,n-\ell\right)$. We remark that (5.4) and (7.1) ensure that
$1\leqslant i_{1}(\ell),\frac{n-1}{2}+1-\ell,n+j_{2}(\ell)-2\ell,n-\ell,n-2\ell<n.$
Therefore, to make use of Lemma 4.3 we first choose $i_{1}(\ell)$ and $j_{2}(\ell)$ such that
$1\leqslant i_{1}(\ell)<\frac{n-1}{2}+1-\ell<n+j_{2}(\ell)-2\ell<n-\ell,$
which in turn follows from
$1\leqslant i_{1}(\ell)<\frac{n-1}{2}+1-N_{2}\quad\textnormal{and}\quad 1\leqslant j_{2}(\ell)<N_{1}.$ (7.3)
We note that the number of choices for $i_{1}(\ell)$ and $j_{2}(\ell)$ satisfying these conditions are
$\frac{n-1}{2}-N_{2}\quad\textnormal{and}\quad N_{1}-1,$ (7.4)
respectively. It is also clear that (7.3) implies (7.1). Finally, to apply Lemma 4.3 we need to ensure
$n-2\ell\in\left[\frac{n-1}{2}+1-\ell+1,n+j_{2}(\ell)-2\ell-1\right];$
this condition follows from
$\ell\leqslant\frac{n-1}{2}-1\quad\textnormal{and}\quad 1\leqslant j_{2}(\ell).$
In particular, both of these conditions are satisfied for all
$N_{1}\leqslant\ell\leqslant N_{2}<\frac{n}{2}-1$
under the assumption that (7.3) holds. Therefore, as long as (5.4) and (7.3) are satisfied, it follows from Lemma 4.3 that
$x_{i}=y_{i}=0$
for all $1\leqslant i\leqslant n$, which in particular implies that the vectors of $\mathfrak{B}_{\ell}$ form a linearly independent set.
We shall prove that we may choose $i_{1}(\ell)$ and $j_{2}(\ell)$, for each $N_{1}\leqslant\ell\leqslant N_{2}$, such that
$\mathfrak{B}_{\ell}\cap\mathscr{S}=\varnothing,\quad\dim\textnormal{Span}_{\mathbb{R}}\mathfrak{B}_{\ell}=2n-1\quad\textnormal{and}\quad\mathfrak{B}_{\ell^{\prime}}\cap\mathfrak{B}_{\ell}=\varnothing,$ (7.5)
if $N_{1}\leqslant\ell^{\prime}<\ell$. To this end, we define
$\mathfrak{R}_{\ell}=\mathscr{S}\cup\mathfrak{B}_{N_{1}}\cup\mathfrak{B}_{N_{1}+1}\cup\cdots\cup\mathfrak{B}_{\ell-1}\cup\widetilde{\mathfrak{B}}_{\ell}\cup\cdots\cup\widetilde{\mathfrak{B}}_{N_{2}},$
for each $N_{1}\leqslant\ell\leqslant N_{2}$. Since every $\widetilde{\mathfrak{B}}_{\ell^{\prime}}$ contains at most two vectors from each block and each column (recall these terminologies from Definition 5.1), it is clear that $\mathfrak{R}_{N_{1}}$ contains $\leqslant(2(N_{2}-N_{1}+1)+2)$ vectors from each block and each column.
From $\mathfrak{R}_{\ell}$ we remove $\mathbf{c}_{\frac{n-1}{2}+1-\ell,\frac{n-1}{2}+1+\ell}$ and $\mathbf{c}_{n+1-\ell,\ell}$ and add $\mathbf{c}_{i_{1}(\ell),\frac{n-1}{2}+1+\ell}$ and $\mathbf{c}_{n+1-\ell,j_{2}(\ell)}$ to obtain $\mathfrak{R}_{\ell+1}$; therefore, the number of vectors from each block increases at most by one, and so as that for each column. Thus every $\mathfrak{R}_{\ell^{\prime}}$ contains
$\leqslant(2(N_{2}-N_{1}+1)+2+N_{2}-N_{1})=3(N_{2}-N_{1})+4$ (7.6)
vectors from each block and from each column. We shall assume that the following holds
$\frac{n-1}{2}-N_{2}\geqslant N_{1}-1\quad\textnormal{and}\quad N_{1}-3(N_{2}-N_{1})\geqslant 6.$ (7.7)
We note that, as explained above, by choosing $i_{1}(\ell)$ and $j_{2}(\ell)$ to satisfy (7.3) we have $\dim\textnormal{Span}_{\mathbb{R}}\mathfrak{B}_{\ell}=2n-1$. We now show by induction that we may choose $i_{1}(\ell)$ and $j_{2}(\ell)$ so that
$(\mathfrak{B}_{\ell}\setminus\widetilde{\mathfrak{B}}_{\ell})\cap\mathfrak{R}_{\ell}=\varnothing\quad\textnormal{and}\quad\dim\textnormal{Span}_{\mathbb{R}}\mathfrak{B}_{\ell}=2n-1$ (7.8)
hold for each $N_{1}\leqslant\ell\leqslant N_{2}$. We now record the following as a lemma; we omit the proof as it is identical to that of Lemma 6.1.
Lemma 7.1.
Suppose (7.8) holds. Then we have (7.5).
We begin with the base case $\ell=N_{1}$. In this case, we know from (7.4), (7.6) and (7.7) that there are
$\geqslant N_{1}-1-(3(N_{2}-N_{1})+4)\geqslant 1$
choices for both $i_{1}(N_{1})$ and $j_{2}(N_{1})$. To see this, first we recall that $i_{1}(N_{1})$ corresponds to the vector $\mathbf{c}_{i_{1}(N_{1}),\frac{n-1}{2}+1+N_{1}}$ (resp. $\mathbf{c}_{n+1-N_{1},j_{2}(N_{1})}$ for $j_{2}(N_{1})$). We may select any vector of the $N_{1}-1$ vectors satisfying (7.3) from the $(\frac{n-1}{2}+1+N_{1})$-th column (resp. $(n+1-N_{1})$-th block) that does not belong to $\mathfrak{R}_{N_{1}}$. Since there are $\leqslant(2(N_{2}-N_{1}+1)+2+N_{2}-N_{1})$ vectors in $\mathfrak{R}_{N_{1}}$ from the $(\frac{n-1}{2}+1+N_{1})$-th column (resp. $(n+1-N_{1})$-th block), we obtain the above inequality. Therefore, we may select appropriate $i_{1}(N_{1})$ and $j_{2}(N_{1})$ such that
$(\mathfrak{B}_{N_{1}}\setminus\widetilde{\mathfrak{B}}_{N_{1}})\cap\mathfrak{R}_{N_{1}}=\varnothing\quad\textnormal{and}\quad\dim\textnormal{Span}_{\mathbb{R}}\mathfrak{B}_{N_{1}}=2n-1,$
and fix these choices.
Now suppose the statement holds for all values greater than or equal to $N_{1}$ and less than $\ell$, for some $N_{1}<\ell\leqslant N_{2}$. We need to choose $i_{1}(\ell)$ and $j_{2}(\ell)$ such that the desired conditions are satisfied, where we recall that these correspond to the vectors $\mathbf{c}_{i_{1}(\ell),\frac{n-1}{2}+1+\ell}$ and $\mathbf{c}_{n+1-\ell,j_{2}(\ell)}$ respectively. These two vectors are to be chosen from the complement of $\mathfrak{R}_{\ell}$. Therefore, by the same argument as in the base case above, we again have
$\geqslant N_{1}-1-(3(N_{2}-N_{1})+4)\geqslant 1$
choices for both $i_{1}(\ell)$ and $j_{2}(\ell)$ such that
$(\mathfrak{B}_{\ell}\setminus\widetilde{\mathfrak{B}}_{\ell})\cap\mathfrak{R}_{\ell}=\varnothing,\quad\textnormal{and}\quad\dim\textnormal{Span}_{\mathbb{R}}\mathfrak{B}_{\ell}=2n-1$
holds. This completes our induction. In conclusion, we have obtained pairwise disjoint $\mathfrak{B}_{N_{1}},\ldots,\mathfrak{B}_{N_{2}}$ satisfying
$\mathfrak{B}_{\ell}\cap\mathscr{S}=\varnothing\quad\textnormal{and}\quad\dim\textnormal{Span}_{\mathbb{R}}\mathfrak{B}_{\ell}=2n-1,$
for each $N_{1}\leqslant\ell\leqslant N_{2}$, provided (5.4) and (7.7) hold.
7.1. Final bound for $T$
By the same argument as in Section 6.1, it follows that
$T\geqslant N_{2}-N_{1}+1.$
Since the only restrictions we needed were (5.4) and (7.7), we set
$N_{1}=\frac{n+1}{2}-N_{2}\quad\textnormal{and}\quad N_{2}=\left\lfloor\frac{2n-4}{7}\right\rfloor<\frac{n}{2}-1$
so that they are both satisfied. Then the above bound for $T$ becomes
$T\geqslant 2N_{2}-\frac{n+1}{2}+1\geqslant 2\left(\frac{2n-4}{7}-1\right)-\frac{n+1}{2}+1=\frac{n-37}{14}.$
Appendix A Theorem 2.2 for small values of $n$ (by Diyuan Wu)
A.1. Introduction
In this appendix, we denote by $M_{0}(n)=M_{0}$ the $(2n+1)\times n^{2}$ coefficient matrix defined in (5.1). Our goal is to prove $T_{0}\geqslant 7$ with the notation in Section 2, that is to find at least $7$ disjoint invertible $(2n+1)\times(2n+1)$ submatrices of $M_{0}(n)$, for $50\leqslant n\leqslant 140$.
In order to define the algorithm we use in the next section, we introduce a few notations:
  • We denote by $\mathbf{c}_{i,j}$ the $((i-1)n+j)$-th column of $M_{0}(n)$ for $1\leqslant i,j\leqslant n$.
  • We call the set of columns $\{\mathbf{c}_{i,1},\ldots,\mathbf{c}_{i,n}\}$ the $i$-th block for each $1\leqslant i\leqslant n$.
  • We define
    $\mathscr{S}_{1}=\{\mathbf{c}_{1,1},\mathbf{c}_{2,2},\ldots,\mathbf{c}_{n,n}\}\quad\textnormal{and}\quad\mathscr{S}_{2}=\{\mathbf{c}_{1,n},\mathbf{c}_{2,n-1},\ldots,\mathbf{c}_{n,1}\}.$
The columns in $\mathscr{S}_{1}$ are precisely the columns of $M_{0}(n)$ with $1$ in the $2n$-th entry, and the columns in $\mathscr{S}_{2}$ are the ones with $1$ in the $(2n+1)$-th entry. We remark that while $\mathscr{S}_{1}$ is as defined in (5.2), $\mathscr{S}_{2}$ is not exactly as in (5.3) since we don’t remove a column from $\mathscr{S}_{1}$; however, this modification is taken care of in the algorithm.
A.2. Algorithm
We now describe the algorithm to find disjoint invertible $(2n+1)\times(2n+1)$ submatrices. The algorithm consists of two parts: the algorithm findIndpSubmatrix (Algorithm 1) to find disjoint invertible submatrices for a given range $(N_{1},N_{2})$ and the procedure Main (Algorithm 2) to enumerate all possible ranges $(N_{1},N_{2})$.
The algorithm findIndpSubmatrix implements the approach described in Sections 57. We briefly summarize the procedure here for completeness. For each $N_{1}<\ell<N_{2}$, we first select a set $\widetilde{\mathfrak{B}}_{\ell}$ of $12n-$ 1 columns containing two columns from each block except the $n$-th column of the $(n-2\ell)$-th block. By the definition of $M_{0}(n)$, there is precisely one column in $\widetilde{\mathfrak{B}}_{\ell}$ with $1$ in the $2n$-th entry and one column with $1$ in the $(2n+1)$-th entry. We remove these two columns and add two columns with 00 in the last two entries. For this purpose, we choose one vector from each of the following two sets:
  • $\{\mathbf{c}_{\frac{n}{2}+1-\ell,j_{1}}:2N_{2}+1\leqslant j_{1}\leqslant\frac{n}{2}+N_{1}\}$ and $\{\mathbf{c}_{n+1-\ell,j_{2}}:1\leqslant j_{2}\leqslant N_{1}\}$ if $n$ is even.
  • $\{\mathbf{c}_{i_{1},\frac{n-1}{2}+1+\ell}:1\leqslant i_{1}\leqslant\frac{n-1}{2}+1-N_{2}\}$ and $\{\mathbf{c}_{n+1-\ell,j_{2}}:1\leqslant j_{2}\leqslant N_{1}\}$ if $n$ is odd.
We enumerate all the possible choices. We call the resulting set of columns $\mathfrak{B}_{\ell}$. Note that since $\mathfrak{B}_{\ell}$ consists of $12n-$ 1 columns with 00 in the last two entries, if the columns of $\mathfrak{B}_{\ell}$ are linearly independent, then the linear independency is preserved under the addition of one column from each $\mathscr{S}_{1}$ and $\mathscr{S}_{2}$. Therefore, we obtain an invertible $(2n+1)\times(2n+1)$ submatrix of $M_{0}(n)$ by considering the submatrix formed by these $12n+$ 1 columns. Finally, to guarantee the disjointness, once we find a linear independent set of $12n+$ 1 columns, we set these columns to be the zero vector in $M_{0}(n)$. Thus, in the later process, any set that contains a column already used will be linearly dependent.
In the procedure Main (Algorithm 2), we enumerate $(N_{1},N_{2})$ to find as many disjoint invertible submatrices as possible:
  • $1\leqslant N_{1}\leqslant\frac{n}{2}$ and $N_{2}=\lfloor\frac{n}{4}\rfloor+\lfloor\frac{N_{1}}{2}\rfloor-1$ if $n$ is even.
  • $2\leqslant N_{2}\leqslant\frac{n-1}{2}$ and $N_{1}=\lfloor\frac{N_{2}}{2}\rfloor$ if $n$ is odd and $n\geqslant 87$ (option 1 in Algorithm 2).
  • $2\leqslant N_{2}\leqslant\frac{n-1}{2}$ and $1\leqslant N_{1}\leqslant N_{2}$ if $n$ is odd and $n<87$ (option 2 in Algorithm 2).
Note that two separate cases are considered when $n$ is odd to save computation time.
Input $n,N_{1},N_{2},M_{0}(n)$;
Initialize $\mathfrak{B}=\varnothing$;
for $\ell=N_{1},\ldots,N_{2}$ do
       Select the following set of columns:
$\widetilde{\mathfrak{B}}_{\ell}=\{\mathbf{c}_{1,1+2\ell-1},\mathbf{c}_{2,2+2\ell-1},\ldots,\mathbf{c}_{n,n+2\ell-1}\}\bigcup\left(\{\mathbf{c}_{1,1+2\ell},\mathbf{c}_{2,2+2\ell},\ldots,\mathbf{c}_{n,n+2\ell}\}\setminus\{\mathbf{c}_{n-2\ell,n}\}\right).$
      if $n$ is even then
             for $j_{1}=2N_{2}+1,\ldots,\frac{n}{2}+N_{1}$ do
                   for $j_{2}=1,\ldots,N_{1}$ do
                         Select the following set of columns:
$\mathfrak{B}_{\ell}=(\widetilde{\mathfrak{B}}_{\ell}\setminus\{\mathbf{c}_{\frac{n}{2}+1-\ell,\frac{n}{2}+\ell},\mathbf{c}_{n+1-\ell,\ell}\})\bigcup\{\mathbf{c}_{\frac{n}{2}+1-\ell,j_{1}},\mathbf{c}_{n+1-\ell,j_{2}}\}.$
if $\mathfrak{B}_{\ell}$ is linearly independent then
                               Randomly select $\mathbf{c}_{1}\in\mathscr{I}_{1}$ and $\mathbf{c}_{2}\in\mathscr{I}_{2}$;
                               $\mathfrak{B}_{\ell}=\mathfrak{B}_{\ell}\bigcup\{\mathbf{c}_{1},\mathbf{c}_{2}\}$;
                               if $\mathfrak{B}_{\ell}$ is linearly independent then
                                    $\mathfrak{B}=\mathfrak{B}\bigcup\mathfrak{B}_{\ell}$;
                               end if
                              Set all the columns in $\mathfrak{B}_{\ell}$ to be the zero vector.
                         end if
                   end for
             end for
       end if
      if $n$ is odd then
             for $i_{1}=1,\ldots,\frac{n-1}{2}+1-N_{2}$ do
                   for $j_{2}=1,\ldots,N_{1}$ do
                         Select the following set of columns:
$\mathfrak{B}_{\ell}=(\widetilde{\mathfrak{B}}_{\ell}\setminus\{\mathbf{c}_{\frac{n-1}{2}+1-\ell,\frac{n-1}{2}+1+\ell},\mathbf{c}_{n+1-\ell,\ell}\})\bigcup\{\mathbf{c}_{i_{1},\frac{n-1}{2}+1+\ell},\mathbf{c}_{n+1-\ell,j_{2}}\}.$
if $\mathfrak{B}_{\ell}$ is linearly independent then
                               Randomly select $\mathbf{c}_{1}\in\mathscr{I}_{1}$ and $\mathbf{c}_{2}\in\mathscr{I}_{2}$;
                               $\mathfrak{B}_{\ell}=\mathfrak{B}_{\ell}\bigcup\{\mathbf{c}_{1},\mathbf{c}_{2}\}$;
                               if $\mathfrak{B}_{\ell}$ is linearly independent then
                                    $\mathfrak{B}=\mathfrak{B}\bigcup\mathfrak{B}_{\ell}$;
                               end if
                              Set all the columns in $\mathfrak{B}_{\ell}$ to be the zero vector.
                         end if
                   end for
             end for
       end if
end for
return $M_{0}(n),\mathfrak{B}$
Algorithm 1 findIndpSubmatrix
Input $n$;
Initialize $\mathfrak{B}=\varnothing$;
if $n$ is even then
       for $N_{1}=1,\ldots,\frac{n}{2}$ do
             $N_{2}=\lfloor\frac{n}{4}\rfloor+\lfloor\frac{N_{1}}{2}\rfloor-1$;
             $M_{0}(n),\mathfrak{B}_{N_{1},N_{2}}=\texttt{findIndpSubmatrix}(n,N_{1},N_{2},M_{0}(n))$;
             $\mathfrak{B}=\mathfrak{B}\bigcup\mathfrak{B}_{N_{1},N_{2}}$;
       end for
end if
if $n$ is odd then
       for $N_{2}=2,\ldots,\frac{n-1}{2}$ do
[…]
             $M_{0}(n),\mathfrak{B}_{N_{1},N_{2}}=\texttt{findIndpSubmatrix}(n,N_{1},N_{2},M_{0}(n))$;
             $\mathfrak{B}=\mathfrak{B}\bigcup\mathfrak{B}_{N_{1},N_{2}}$;
       end for
end if
return $\mathfrak{B}$
Algorithm 2 Main
A.3. Results and verification
Both the algorithm and the results can be found in the following public GitHub repository: https://github.com/DiyuanWu/find_indpsubmatrix.git.
In Table 1 below, we record the number of disjoint invertible $(2n+1)\times(2n+1)$ submatrices of $M_{0}(n)$ found by the algorithm. It can be seen that we can find at least $7$ disjoint invertible submatrices, for each $50\leqslant n\leqslant 140$, as desired. We also store the submatrices found for each $n$ in the file result_dict; this is done by storing the indices of the columns forming each submatrix (note that the indices start from 00 in Python).
We also provide the verification of the results. In particular, we need to verify: (1) each submatrix found is invertible, and (2) all the submatrices are disjoint. To verify the invertibility, we check whether the submatrix is of full rank. To verify the disjointness, we first aggregate all the columns of the submatrices found into one set using the set item in Python, which in particular does not allow any duplicates, and then we check whether the number of the columns in this set is as expected. The results can be verified by running the file verify.py.
$n$ number of disjoint invertible submatrices $n$ number of disjoint invertible submatrices
50 12 51 13
52 15 53 13
54 14 55 14
56 16 57 14
58 17 59 14
60 17 61 15
62 18 63 15
64 19 65 16
66 19 67 17
68 19 69 17
70 20 71 18
72 21 73 18
74 22 75 19
76 23 77 19
78 23 79 20
80 25 81 21
82 24 83 22
84 26 85 22
86 27 87 23
88 27 89 24
90 28 91 25
92 27 93 25
94 30 95 26
96 30 97 27
98 30 99 28
100 30 101 28
102 31 103 29
104 31 105 29
106 32 107 30
108 34 109 30
110 35 111 31
112 36 113 32
114 36 115 33
116 38 117 33
118 37 119 34
120 38 121 35
122 38 123 35
124 40 125 36
126 40 127 36
128 41 129 36
130 42 131 38
132 43 133 39
134 43 135 40
136 44 137 40
138 45 139 41
140 44
Table 1. Number of disjoint invertible submatrices found
Email address: Diyuan.Wu@ist.ac.at
IST Austria, Am Campus 1, 3400 Klosterneuburg, Austria.